OpenStudy (anonymous):

solve -3(2x-1)=x+10. show all your work

OpenStudy (anonymous):

-6x+3=x+10 -7x=7 x=-1

OpenStudy (anonymous):

\[ x= -1\]

OpenStudy (anonymous):

i can tell u if u did n't understand

OpenStudy (anonymous):

do u want me too answer this

OpenStudy (anonymous):

it's ok, thank you man, i got it